# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v5

@@ Abstract
To explain how dissipative active turbulence sustains inertial cascades, we construct a chiral model mapping overdamped agents to supersonic topological gas dynamics. Here, Mach cones function as acoustic horizons, shielding defect cores from the sound radiation of shallow water flows. We show that disorder activates a topological heat pump driving an inverse cascade toward a negative-temperature Onsager dipole, unless arrested into a vortex glass by insufficient dispersion. This mapping identifies a universality class unifying active swarms with the statistical mechanics of classical inviscid fluids.
@@ I Introduction
Though active matter systems frequently and demonstrably exhibit viscous, dissipative turbulence [14, 15, 16], the observational discrepancy implies a missing link in the theoretical framework. While odd viscosity enables unidirectional flow [17, 18] and recent models suggest swarming can be entropy-driven [19], the thermodynamic end-state of such systems remains largely unexplored. Specifically, does the breaking of detailed balance [20] by intrinsic rotation sustain a chaotic steady state, or can it drive the system toward a specific macroscopic attractor?
In this letter, we propose that the answer to the conundrum lies in treating such systems as topological gases. By simulating chiral active matter as an ensemble of $N$ locally phase-coupled oscillators experiencing a Kuramoto-like phase-interaction [21], we demonstrate that active turbulence can exhibit a duality; a microscopic active bath dominated by enstrophy injection can simultaneously be described as a macroscopic renormalised fluid governed by the conservative advection of topological defects. By filtering out the microscopic singularities, we reveal that the simulated system undergoes a “hidden” inverse energy cascade, indistinguishable from Euler (frictionless, inviscid) turbulence. This cascade functions akin to a thermodynamic pump, driving the system toward an Onsager dipole, a negative-temperature condensate of vorticity [22], thereby unifying the phenomenology of biological swarms with the rigorous statistical mechanics of point vortices.
@@ II Methodology
We define the system as a chiral active fluid composed of $N=50,000$ particles, or discrete agents, in a bounded periodic domain $\mathcal{D}\in\mathbb{R}^{2}$. The state of the $i^{\text{th}}$ agent is defined by its position $\mathbf{r}_{i}(t)$ and its internal phase $\phi_{i}(t)\in[0,2\pi)$. The agents do not interact via pairwise collisions (as in granular matter) but through a local mean field. Following Ref. [21], we define the complex order parameter field $Z(\mathbf{r},t)$ as the convolution of the microscopic agent distribution with a finite-range interaction kernel $G$:
@@
Following Ref. [21], the internal phase $\phi_{i}$ evolves according to an overdamped, localized, driven Kuramoto-Sakaguchi-like interaction [23, 24]:
@@
where $\omega_{i}$ are natural frequencies drawn from a power-law distribution $P(\omega)\sim\omega^{-n}$, $a_{0}$ is the coupling Strength, $\eta_{i}$ is a Gaussian white noise term (amplitude 0.15), and $n=1.5$ is a positive constant of order unity to facilitate the non-linear transfer of energy (see Refs. [21, 25]). $R_{i}$, the local order parameter, ensures that synchronisation is triggered by initially random local alignments in phase. $\omega_{i}$ (and the stochastic term $\eta_{i}$) prevent a “ferromagnetic collapse”, in which the system simply synchronises globally ($\phi_{i}=\text{const}$) causing a uniform ballistic drift. By forcing the particles to oscillate with an inherent frequency (or attempt to do so), the $\omega_{i}$ terms act as frustration, constantly injecting vorticity at the micro-scale, which, as we shall demonstrate, feeds an inverse energy cascade.
@@
At this point, We identify the characteristic scales in out model; the kernel size $\sigma$ and the active swim speed $v_{0}$, which yield the interaction time $\tau\equiv\sigma/v_{0}=(3\text{ grid units})/(0.5\text{ grid units/s})=6$ s, yielding the dimensionless units $\tilde{x}=x/\sigma$, $\tilde{t}=t/\tau$, and $\tilde{\omega}=\omega(\sigma/v_{0})$ (we subsequently drop the tilde). The system size becomes $\Lambda=2L/\sigma\approx 13$. A natural definition of the RFE is then afforded by $\tau$; we define the RFE as the center of mass of a particle’s trajectory during $\tau$, effectively course-graining the singularities:
@@
Figure 1: The emergence of the Onsager dipole in simulated chiral active matter, shown in seven temporal snapshots of the simulation space. The top row shows the RFE particle point-clouds $r_{i}^{\text{RFE}}$, the middle row shows topological charge $m$, and the bottom row shows inertial mass $R^{2}$, using the colorscales indicated.
@@
By decomposing the particle velocity into a mean flow with fluctuations, we find that the macroscopic velocity $\mathbf{u}$ is governed by the mean phase gradient $\nabla\Psi$ scaled by the local order parameter $R$: $\mathbf{u}\approx R\nabla\Psi$. While we derive a set of constraining equations that are isomorphic to shallow water hydrodynamics, there are striking differences: the effective inertial mass in our model scales with order:
$\lambda=R^{2},$ (6)
with an effective restoring potential (enthalpy),
$h_{\text{eff}}=\frac{v_{0}^{2}}{2}R^{2}.$ (7)
The macroscopic restoring force is governed by the Eq. (7)’s gradient, enforcing a sound speed that is also proportional to order,
$c_{s}=\frac{v_{0}R}{\sqrt{2}}.$ (8)
Since the inertial velocity likewise scales with $R$, this locks the flow into an inherently supersonic state (Mach number >1>1). Consequently, defect cores ($R\rightarrow 0$) are isolated behind Mach cones, creating sonic horizons that prevent radiative decay. While fast compressible shocks disperse inertial energy locally, the time-averaged background flow remains incompressible. This effectively traps energy within the topological degrees of freedom, meaning the macroscopic dynamics are governed strictly by the conservation of topological charge.
@@ III Results
Finally, we demonstrate that the formation of the dipole is a thermodynamic inevitability. Figure 2d) maps the system’s trajectory in the energy-order phase plane ($E$ vs. $P$). The system climbs away from the disordered origin (random gas), increasing both its macroscopic kinetic energy and its dipole order parameter $P$, while the flucutations in Figure 2d) reflect the dynamic shock merger mechanism. In the context of Eq. (10), we observe an ascent up the entropy curve into the regime of negative absolute temperature [22]. In the active bath, intrinsic disorder ($\omega_{i}$) continuously injects enstrophy to maximize the entropy of the renormalised fluid, forcing the long-term spectral condensation of vorticity as the only accessible maximum-entropy state within the bounded domain. The negative spectral flux in Figure 2b) confirms our hypothesis that the modeled particles act as a dissipative kinetic layer that drives a superfluid governed by the conservative Kirchhoff-Onsager Hamiltonian, unifying the phenomenology of chiral active matter with the rigorous statistical mechanics of classical inviscid fluids.
@@ IV Discussion
The realization of the Onsager dipole depends strictly on the system’s ability to overcome topological energy barriers [25]. As alluded to by Ref. [21], we find that the intrinsic frequency dispersion $\Delta\omega$ determines whether the topological gas freezes, jams, or flows.
@@ V Conclusion
We have identified a disorder-induced transition in chiral active matter where a topological heat pump drives an inverse cascade, saturating in a negative-temperature, dynamic Onsager dipole, representing a hydrodynamic limit that is isomorphic with shallow water theory. This mapping reconciles dissipative microphysics with macroscopic inertial transport. Our findings suggest active matter is governed by an emergent topological inertia, offering a new paradigm for robust transport.
@@ VII.1 Velocity and continuity
for a peaked (narrow) Gaussian distribution in $\delta\phi$; we then drop the subscript. $R$ measures the degree of of local synchronisation. Thus, we obstain,
@@
Eq. (16) is globally valid: at defect cores where the phase $\Psi$ is undefined, the order parameter vanishes ($R\to 0$), ensuring the velocity field remains continuous and bounded ($\mathbf{u}\to 0$). Recognizing that $\nabla\Psi$ scales as $1/r$ near defects, we note that the vector $\mathbf{u}$ is parallel to $\nabla\Psi$. Far from defect cores, where the phase gradient magnitude varies slowly ($|\nabla\Psi|\approx\text{const}$), the flow simplifies to an effective potential flow:
$\mathbf{u}\approx R\nabla\Psi.$ (17)
@@ VII.2 Momentum and Pressure
Matching the anisotropic coefficients (the $\cos(2\Psi)$ terms) determines the coherence coefficient:
@@ VII.3 Shallow waters and the active enthalpy functional
As stated in the Methodology section, the interpretation of Eq. (19) identifies the system being modeled as isomorphic to the shallow water framework, where the active density $\rho$ acts as the fluid height. To proceed, we shall use the active enthalpy functional as a theoretical vehichle [68, 69]. However, we must first bolster the connection with a specific empirical observation. Figure 7 shows the end-state of a Onsager dipole simulation, color-coding $\omega_{i}$ (top), $R$ (middle), and topological charge $m$ (bottom). We make two important observations;
@@
(2) Figure 7a–d) present clear, unambiguous empirical evidence for the fact that, at topological defect cores, $R\to 0$ and $\omega_{i}\gg\bar{\omega}$, validating Eq. (16) and other instances where the mean field $\Psi$ appears in the numerator. As we shall motivate below, the conditions at defect cores implies the fluid is stiff. The flow must immediately adjust to the topography defined by $\omega(\mathbf{r})$. The segregation of the active particles constitutes the a potential $\Phi(\mathbf{r})$ that forces the flow.
Having motivated a conceptual shift from a Lagrangian to an Eulerian perspective, we define the effective fluid density as $\rho(\mathbf{r},t)\equiv R(\mathbf{r},t)^{2}$. We then use $\omega(\mathbf{r})$ as an external scalar potential produced by intrinsic frequency mismatch $\omega_{i}-\bar{\omega}$, $\bar{\omega}$ being mean frequency, the ensemble average of which equals dispersion $\Delta\omega$.
The total potential energy density $\mathcal{U}_{\text{tot}}$ is composed of the internal interaction energy (barotropic fluid self-interaction, or non-linear coupling in turbulence physics), and the external topographic energy [70]. We have,
$\mathcal{U}_{\text{tot}}[\rho]=\underbrace{g_{\text{eff}}/2(\rho-\rho_{0})^{2}}_{\text{Interaction}}+\underbrace{\gamma\rho\cdot\omega(\mathbf{r})}_{\text{Topography}},$ (32)
where $g_{\text{eff}}$ is the effective coupling constant representing the stiffness of the order parameter (analogous to gravity $g$ in shallow waters), and $\gamma$ is a coupling coefficient with dimensions $\text{length}^{2}\text{time}^{-1}$ relating frequency disorder $\omega(\mathbf{r})$ to potential energy. Two concise observations can now be had by dimensional analysis. (1) the units of $\gamma$ naturally lead to the definition $\gamma=\sigma v_{0}$, anchoring the topographic potential to the physical interaction scale of the agents. (2) by direct comparison between Eqs. (LABEL:eq:peff2) and (32), we observe that $g_{\text{eff}}\equiv v_{0}^{2}/2$.
@@ VII.4 Active acoustic speed
At this point, we must address acoustics, as a shallow water vortex diffuses into the medium through sound waves. To demonstrate why our Onsager dipole is unable to do so, we first revisit the continuity and momentum equations (Eqs. 18 and 36) using $\rho=R^{2}$,
@@
by merit of Eqs. (15) and (45). The macroscopic flow is therefore inherently supersonic ($M>1$). Near a defect, however, $R\to 0$, and so $c_{s}\to 0$; state information cannot escape a defect core, turning the latter into sonic black holes. This implies the existence of a “sonic horizon” around defect cores, which become isolated behind Mach cones. Hence, defect cores cannot diffuse into their surroundings by radiating sound (such as geophysical vortices), justifying our appliation of the the point-vortex Onsager-Kirchoff Hamiltonian.

### v6

@@ Abstract
To explain how dissipative active turbulence sustains inertial cascades, we map overdamped chiral flocks to a compressible vortex gas, realizing supersonic topological gas dynamics. Here, Mach cones function as acoustic horizons, shielding defect cores from the radiative decay of shallow water flows. Disorder activates a topological heat pump driving an inverse cascade toward a negative-temperature Onsager dipole, unless arrested into a vortex glass. This identifies a universality class unifying active swarms with classical inviscid fluids.
@@ I Introduction
Though active matter systems frequently and demonstrably exhibit viscous, dissipative turbulence [14, 15, 16], the observational discrepancy implies a missing link in the theoretical framework. While odd viscosity enables unidirectional flow [17, 18] and recent models suggest swarming can be entropy-driven [19], the thermodynamic end-state of such systems remains largely unexplored. Specifically, does the breaking of detailed balance [20] by intrinsic rotation lead to a chaotic steady state, or can it drive the system toward a specific macroscopic attractor?
In this letter, we propose that the answer to the conundrum lies in treating such systems as topological gases. By simulating chiral active matter as an ensemble of $N$ locally phase-coupled oscillators experiencing a Kuramoto-like phase-interaction [21, 22], we demonstrate that active turbulence can exhibit a duality; a microscopic active bath dominated by enstrophy injection can simultaneously be described as a macroscopic renormalised fluid governed by the conservative advection of topological defects. By filtering out the microscopic singularities, we reveal that the simulated system undergoes a “hidden” inverse energy cascade, indistinguishable from Euler (frictionless, inviscid) turbulence. This cascade functions akin to a thermodynamic pump, driving the system toward an Onsager dipole, a negative-temperature condensate of vorticity [23], thereby unifying the phenomenology of biological swarms with the rigorous statistical mechanics of point vortices.
@@ II Methodology
We define the system as a chiral active fluid composed of $N=50,000$ particles, or discrete agents, in a bounded periodic domain $\mathcal{D}\in\mathbb{R}^{2}$. The state of the $i^{\text{th}}$ agent is defined by its position $\mathbf{r}_{i}(t)$ and its internal phase $\phi_{i}(t)\in[0,2\pi)$. The agents do not interact via pairwise collisions (as in granular matter) but through a local mean field. Following Refs. [21], we define the complex order parameter field $Z(\mathbf{r},t)$ as the convolution of the microscopic agent distribution with a finite-range interaction kernel $G$:
@@
Following Ref. [21], the internal phase $\phi_{i}$ evolves according to an overdamped, localized, driven Kuramoto-Sakaguchi-like interaction [24], recently identified as strictly isomorphic to the Adler equation [25, 26] for disordered Josephson junction arrays [22],
@@
where $\omega_{i}$ are natural frequencies drawn from a power-law distribution $P(\omega)\sim\omega^{-n}$, $a_{0}$ is the coupling Strength, $\eta_{i}$ is a Gaussian white noise term (amplitude 0.15), and $n=1.5$ is a positive constant of order unity, see Ref. [22] for a study of the role of this parameter). $R_{i}$, the local order parameter, ensures that synchronisation is triggered by initial random local alignments in phase. $\omega_{i}$ (and the stochastic term $\eta_{i}$) prevent a “ferromagnetic collapse”, in which the system simply synchronises globally ($\phi_{i}=\text{const}$) causing a uniform ballistic drift. By forcing the particles to oscillate with an inherent frequency (or attempt to do so), the $\omega_{i}$ terms act as frustration, constantly injecting vorticity at the micro-scale, which, as we shall demonstrate, feeds an inverse energy cascade.
@@
At this point, We identify the characteristic scales in out model; the kernel size $\sigma$ and the active swim speed $v_{0}$, which yield the interaction time $\tau\equiv\sigma/v_{0}=(3\text{ grid units})/(0.5\text{ grid units/s})=6$ s, yielding the dimensionless units $\tilde{x}=x/\sigma$, $\tilde{t}=t/\tau$, and $\tilde{\omega}=\omega(\sigma/v_{0})$ (we subsequently drop the tilde). The system size becomes $\Lambda=2L/\sigma\approx 13$. A natural definition of the RFE is then afforded by $\tau$; we define the RFE as the center of mass of a particle’s trajectory during $\tau$, effectively coarse-graining the singularities:
@@
Figure 1: The emergence of the Onsager dipole in simulated chiral active matter, shown in seven temporal snapshots of the simulation space. The top row shows the raw particle point-clouds, with black denoting ‘trapped’ (superconducting) agents and red denoting ‘running’ (phase slipping) agents [22], the middle row shows topological charge $m$, and the bottom row shows inertial mass $R^{2}$ (with flow lines overlaid), using the colorscales indicated.
@@
By decomposing the particle velocity into a mean flow with fluctuations, we find that the macroscopic velocity $\mathbf{u}$ is governed by the mean phase gradient $\nabla\Psi$ scaled by the local order parameter $R$: $\mathbf{u}\sim R\nabla\Psi$. While we derive a set of constraining equations that are isomorphic to shallow water hydrodynamics, there are striking differences: the effective inertial mass in our model scales with order, $\lambda=R^{2}$, a relationship strictly predicted by the macroscopic phase stiffness of the agent dynamics as a motile, disordered Josephson array [22] (see Supplementary Material, Eq. 30). Moreover, we identify the active enthalpy $h_{\text{eff}}=v_{0}^{2}R^{2}/2$, which implies a sound speed $c_{s}\propto R$. Since the flow velocity $\mathbf{u}$ scales identically, the system is locked into a global supersonic state ($M\approx\sqrt{2}$). Consequently, defect cores are shielded by sonic horizons, preventing radiative decay and enforcing strict conservation of topological charge. In the Adler formalism derived by Ref. [22], these cores correspond to agents that are continuously phase-slipping, unable to sustain a phase supercurrent.
@@ III Results
Finally, we demonstrate that the formation of the dipole is a thermodynamic inevitability. Figure 2d) maps the system’s trajectory in the energy-order phase plane ($E$ vs. $P$). The system climbs away from the disordered origin (random gas), increasing both its macroscopic kinetic energy and its dipole order parameter $P$, while the fluctuations in Figure 2d) reflect the dynamic shock merger mechanism. In the context of Eq. (7), we observe an ascent up the entropy curve into the regime of negative absolute temperature [23]. In the active bath, intrinsic disorder ($\omega_{i}$) continuously injects enstrophy to maximize the entropy of the renormalised fluid, forcing the system deeper into the tilted washboard potential identified in the Adler formalism [22]. This necessitates the long-term spectral condensation of vorticity as the only accessible maximum-entropy state within the bounded domain. The negative spectral flux in Figure 2b) confirms our hypothesis that the modeled particles act as a dissipative kinetic layer that drives a superfluid governed by the conservative Kirchhoff-Onsager Hamiltonian, unifying the phenomenology of chiral active matter with the rigorous statistical mechanics of classical inviscid fluids.
@@ IV Discussion
The realization of the Onsager dipole depends strictly on the system’s ability to overcome topological energy barriers [53]. As in Ref. [21], we find that the intrinsic frequency dispersion $\Delta\omega$ determines whether the topological gas freezes, jams, or flows, reflecting the depinning transition of the tilted washboard potential experienced by each agent, where $\Delta\omega$ acts as the effective bias current driving the system across the Adler-Ohmic bifurcation [22].
@@ V Conclusion
We have identified a disorder-induced transition in chiral active matter where a topological heat pump drives an inverse cascade, saturating in a negative-temperature, dynamic Onsager dipole, representing a hydrodynamic limit that is isomorphic with shallow water theory. This mapping reconciles dissipative microphysics with macroscopic inertial transport. Our findings suggest chiral active matter is governed by an emergent topological inertia, offering a new paradigm for robust transport.
@@ VII.1 Velocity and continuity
for a peaked (narrow) Gaussian distribution in $\delta\phi$; we then drop the subscript. $R$ measures the degree of of local synchronisation. Thus, we obtain,
@@
Eq. (13) is globally valid: at defect cores where the phase $\Psi$ is undefined, the order parameter vanishes ($R\to 0$), ensuring the velocity field remains continuous and bounded ($\mathbf{u}\to 0$). Recognizing that $\nabla\Psi$ scales as $1/r$ near defects, we note that the vector $\mathbf{u}$ is parallel to $\nabla\Psi$. Far from defect cores, where the phase gradient magnitude varies slowly ($|\nabla\Psi|\approx\text{const}$), the flow simplifies to an effective potential flow. Far from defect cores, where the phase gradient varies slowly, the flow simplifies to an effective potential flow. By identifying the effective circulation quantum $\kappa=v_{0}\sigma$ (see Eq. 29), demanded by the phase stiffness of the agent description as a motile, disordered Josephson junction array [22], we write:
$\mathbf{u}\approx\kappa R\nabla\Psi=v_{0}\sigma R\nabla\Psi.$ (14)
@@ VII.2 Momentum and Pressure
Matching the anisotropic coefficients (the $\cos 2\Psi$ terms) determines the coherence coefficient:
@@
While the foregoing provided a useful and instructional way to explicitly derive the emergent inertial mass through stress-tensor matching, a more rigorous derivation follows from the isomorphism to the Josephson junction array established in Ref.  [22]. In this framework, the macroscopic superconductivity maintains a stiffness torque mediated by the complex order parameter field $Z=Re^{i\Psi}$. Crucially, Ref. [22] identified that the agent dynamics are formally isomorphic to the Adler equation, meaning individual agents move in a tilted washboard potential [69],
$V_{\text{eff}}(\phi_{i})=-a_{0}R\cos(\Psi-\phi_{i})-\omega_{i}\phi_{i}.$ (28)
According to the Ginzburg-Landau formalism for phase transitions, the free energy cost of sustaining a phase supercurrent (the superfluidic stiffness) scales with the squared magnitude of the complex order parameter field [70],
$\mathcal{E}_{kin}=\frac{1}{2}|Z|^{2}(\nabla\Psi)^{2}=\frac{1}{2}R^{2}(\nabla\Psi)^{2}.$ (29)
We identifying the gradient term as the superfluid velocity $v_{s}=\kappa\nabla\Psi$ (where $\kappa$ is a circulation quantum). By comparing this energy density to the hydrodynamic kinetic energy form $\lambda v_{s}^{2}/2$, we immediately recover the identity,
$\lambda\equiv R^{2},$ (30)
confirming that the effective inertial mass $\lambda$ derived in our hydrodynamic limit (Eq. 27) corresponds physically to the superfluid density (or phase stiffness) of the underlying Josephson junction array.
@@ VII.3 Shallow waters and the active enthalpy functional
As stated in the Methodology section, the interpretation of Eq. (16) identifies the system being modeled as isomorphic to the shallow water framework, where the active density $\rho$ acts as the fluid height. To proceed, we shall use the active enthalpy functional as a theoretical vehicle [72, 73]. However, we must first bolster the connection with a specific empirical observation. Figure 7 shows the end-state of a Onsager dipole simulation, color-coding $\omega_{i}$ (top), $R$ (middle), and topological charge $m$ (bottom). We make two important observations;
@@
(2) Figure 7a–d) present clear, unambiguous empirical evidence for the fact that, at topological defect cores, $R\to 0$ and $\omega_{i}\gg\bar{\omega}$, which, along with Eq. (17) validates Eqs. (13). As we shall motivate below, the conditions at defect cores implies the fluid is stiff. The flow must immediately adjust to the topography defined by $\omega(\mathbf{r})$. The segregation of the active particles constitutes the a potential $\Phi(\mathbf{r})$ that forces the flow.
Having motivated a conceptual shift from a Lagrangian to an Eulerian perspective, we define the effective fluid density as $\rho(\mathbf{r},t)\equiv R(\mathbf{r},t)^{2}$. We then use $\omega(\mathbf{r})$ as an external scalar potential produced by intrinsic frequency mismatch $\omega_{i}-\bar{\omega}$, where $\bar{\omega}$ is the mean frequency. The ensemble average of this quantity then equals dispersion $\Delta\omega$.
To derive the macroscopic energy functional $\mathcal{U}_{\text{tot}}$, we distinguish between the active pressure and the disorder potential for individual agents derived in Ref. [22]. First, the effective fluid self-interaction arises from the active stress tensor $\langle v_{i}v_{j}\rangle$ derived in Eq. (26). The isotropic component of this stress necessitates a macroscopic swim pressure that scales quartically with order,
$P_{\text{swim}}\sim v_{0}^{2}R^{4},$ (32)
This Equation of State ($P\propto\rho^{2}$, where $\rho=R^{2}$) corresponds to the barotropic potential energy density of a shallow water fluid.
We define the total energy functional $\mathcal{U}_{\text{tot}}[\rho]$ by combining this barotropic term with the topographic potential:
$\mathcal{U}_{\text{tot}}[\rho]=\frac{g_{eff}}{2}(\rho-\rho_{0})^{2}+\gamma\rho\cdot\omega(r),$ (33)
where $g_{\text{eff}}$ is the effective coupling constant representing the stiffness of the order parameter (analogous to gravity $g$ in shallow waters), and $\gamma$ is a coupling coefficient with dimensions $\text{length}^{2}\text{time}^{-1}$ relating frequency disorder $\omega(\mathbf{r})$ to potential energy. Here, the frequency dispersion $\omega_{i}$, acting as a local conservative potential in the Josephson dynamics of the individual agents [22], maps to the external topographic potential $\omega(r)$ in the continuum limit (Figure 7). We note that Eq. (33) corresponds directly to the potential energy of a fluid flowing over variable topography [74].
Two concise observations can now be made by dimensional analysis. (1) The units of $\gamma$ naturally lead to the definition $\gamma=\sigma v_{0}$, anchoring the topographic potential to the physical interaction scale of the agents. (2) By direct comparison between the divergence of the isotropic stress in Eq. (26) and the shallow water pressure gradient, we observe that $g_{\text{eff}}\equiv v_{0}^{2}/2$. This follows from observing that the active stress force $\nabla\cdot(\rho\frac{v_{0}^{2}}{2})$ corresponds to the shallow water pressure force $g_{\text{eff}}\nabla\rho$.
@@ VII.4 Active acoustic speed
At this point, we must address acoustics, as a shallow water vortex diffuses into the medium through sound waves. To demonstrate why our Onsager dipole is unable to do so, we first revisit the continuity and momentum equations (Eqs. 15 and 37) using $\rho=R^{2}$,
@@
by merit of Eqs. (12) and (46). The macroscopic flow is therefore inherently supersonic ($M>1$). Near a defect, however, $R\to 0$, and so $c_{s}\to 0$; state information cannot escape a defect core, turning the latter into sonic black holes, aply illustrated by the red ‘trapped’ agents being effectively confined to the vortex cores in Figure 1. This implies the existence of a “sonic horizon” around defect cores, which become isolated behind Mach cones. Hence, defect cores cannot diffuse into their surroundings by radiating sound (such as geophysical vortices), justifying our appliation of the the point-vortex Onsager-Kirchoff Hamiltonian.